OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x?

OpenStudy (anonymous):

x = -3 x = -2 x = 4 x = 6

OpenStudy (apoorvk):

The mid point of two points A(x,y0 and B(x',y') is given by M(A,B), where: A = (x + x')/2 B = (y + y')/2 You have the values of x', y, y', A and B sated. Find out 'x'.

OpenStudy (anonymous):

So do i plug in, you mean ?

OpenStudy (apoorvk):

yeah, plug-in!

OpenStudy (anonymous):

So would it be ...-2+(-2)/2 over 1+2/2

OpenStudy (apoorvk):

what's your answer?

OpenStudy (apoorvk):

would be: 2 = (x - 2)/2

OpenStudy (anonymous):

i appreciate your help but this really doesn't make sense to me ..:(